# On the Expressiveness of Visual Prompt Experts

### v5

@@ E.1 Per-task Results for VTAB-1K and FGVC
Table 5: VTAB-1K Natural per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
ViT-B/16 [9] VTAB-1K [82] Natural [7]
@@
Full [29] 68.9 87.7 64.3 97.2 86.9 87.4 38.8 75.88
Linear [29] 63.4 85.0 63.2 97.0 86.3 36.6 51.0 68.93 [1]
Partial-1 [81] 66.8 85.9 62.5 97.3 85.5 37.6 50.6 69.44 [2]
MLP-2 [3] 63.2 84.8 60.5 97.6 85.9 34.1 47.8 67.70 [2]
MLP-3 [3] 63.8 84.7 62.3 97.4 84.7 32.5 49.2 67.80 [2]
MLP-5 [3] 59.3 84.4 59.9 96.1 84.4 30.9 46.8 65.98 [1]
MLP-9 [3] 53.1 80.5 53.9 95.1 82.6 24.4 43.7 61.90 [1]
Sidetune [83] 60.7 60.8 53.6 95.5 66.7 34.9 35.3 58.21 [0]
Bias [62] 72.8 87.0 59.2 97.5 85.3 59.9 51.4 73.30 [3]
Adapter-256 [2] 74.1 86.1 63.2 97.7 87.0 34.6 50.8 70.50 [4]
Adapter-64 [2] 74.2 85.8 62.7 97.6 87.2 36.3 50.9 70.65 [4]
Adapter-8 [2] 74.2 85.7 62.7 97.8 87.2 36.4 50.7 70.67 [4]
VPT-Shallow [31] 77.7 86.9 62.6 97.5 87.3 74.5 51.2 76.81 [4]
@@
Table 6: VTAB-1K Specialized per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
ViT-B/16 [9] VTAB-1K [82] Specialized [4]
@@
Full [29] 79.7 95.7 84.2 73.9 83.36
Linear [29] 78.5 87.5 68.6 74.0 77.16 [1]
Partial-1 [81] 78.6 89.8 72.5 73.3 78.53 [0]
MLP-2 [3] 74.3 88.8 67.1 73.2 75.86 [0]
MLP-3 [3] 77.0 88.0 70.2 56.1 72.83 [0]
MLP-5 [3] 73.7 87.2 64.8 71.5 74.31 [0]
MLP-9 [3] 78.5 83.0 60.2 72.3 73.49 [0]
Sidetune [83] 58.5 87.7 65.2 61.0 68.12 [0]
Bias [62] 78.7 91.6 72.9 69.8 78.25 [0]
Adapter-256 [2] 76.3 88.0 73.1 70.5 76.98 [0]
Adapter-64 [2] 76.3 87.5 73.7 70.9 77.10 [0]
Adapter-8 [2] 76.9 89.2 73.5 71.6 77.80 [0]
VPT-Shallow [31] 78.2 92.0 75.6 72.9 79.66 [0]
@@
Table 7: VTAB-1K Structured per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
ViT-B/16 [9] VTAB-1K [82] Structured [8]
@@
Full [29] 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1 47.64
Linear [29] 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2 26.84 [0]
Partial-1 [81] 41.5 34.3 33.9 61.0 31.3 32.8 16.3 22.4 34.17 [0]
MLP-2 [3] 45.2 31.6 31.8 55.7 30.9 24.6 16.6 23.3 32.47 [0]
MLP-3 [3] 47.8 32.8 32.3 58.1 12.9 21.2 15.2 24.8 30.62 [0]
MLP-5 [3] 50.8 32.3 31.5 56.4 7.5 20.8 14.4 20.4 29.23 [0]
MLP-9 [3] 47.5 27.9 28.9 54.0 6.2 17.7 10.8 16.2 26.15 [0]
Sidetune [83] 27.6 22.6 31.3 51.7 8.2 14.4 9.8 21.8 23.41 [0]
Bias [62] 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1 44.09 [2]
Adapter-256 [2] 45.7 37.4 31.2 53.2 30.3 25.4 13.8 22.1 32.39 [0]
Adapter-64 [2] 42.9 39.9 30.4 54.5 31.9 25.6 13.5 21.4 32.51 [0]
Adapter-8 [2] 45.2 41.8 31.1 56.4 30.4 24.6 13.2 22.0 33.09 [0]
VPT-Shallow [31] 50.5 58.6 40.5 67.1 68.7 36.1 20.2 34.1 46.98 [4]
@@
Table 8: FGVC per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in [$\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
ViT-B/16 [9] FGVC [5]
@@
Full [29] 87.3 82.7 98.8 89.4 84.5 88.54
Linear [29] 85.3 75.9 97.9 86.2 51.3 79.32 [0]
Partial-1 [81] 85.6 77.8 98.2 85.5 66.2 82.63 [0]
MLP-2 [3] 85.7 77.2 98.2 85.4 54.9 80.28 [0]
MLP-3 [3] 85.1 77.3 97.9 84.9 53.8 79.80 [0]
MLP-5 [3] 84.2 76.7 97.6 84.8 50.2 78.71 [0]
MLP-9 [3] 83.2 76.0 96.2 83.7 47.6 77.31 [0]
Sidetune [83] 84.7 75.8 96.9 85.8 48.6 78.35 [0]
Bias [62] 88.4 84.2 98.8 91.2 79.4 88.41 [3]
Adapter-256 [2] 87.2 84.3 98.5 89.9 68.6 85.70 [2]
Adapter-64 [2] 87.1 84.3 98.5 89.8 68.6 85.67 [2]
Adapter-8 [2] 87.3 84.3 98.4 88.8 68.4 85.46 [1]
VPT-Shallow [31] 86.7 78.8 98.4 90.7 68.7 84.62 [1]
@@ E.2 Per-task Results on MAE and MoCo v3
Table 9: VTAB-1K Natural per-task results for ViT-B/16 pre-trained on MAE [21]. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-B/16 [9] VTAB-1K [82] Natural [7]
@@
Table 10: VTAB-1K Specialized Per-Task Results for ViT-B/16 Pre-trained on MAE [21]. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-B/16 [9] VTAB-1K [82] Specialized [4]
@@
Table 11: VTAB-1K Structured Per-Task Results for ViT-B/16 Pre-trained on MAE [21]. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-Base/16 [9] VTAB-1K [82] Structured [8]
@@
Table 12: VTAB-1K Natural Per-Task Results for ViT-B/16 Pre-trained on MoCo v3 [4]. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-B/16 [9] VTAB-1K [82] Natural [7]
@@
Table 13: VTAB-1K Specialized Per-Task Results for ViT-B/16 Pre-trained on MoCo v3 [4]. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-B/16 [9] VTAB-1K [82] Specialized [4]
@@
Table 14: VTAB-1K Structured Per-Task Results for ViT-B/16 Pre-trained on MoCo v3 [4]. “Number of Wins” in [$\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-Base/16 [9] VTAB-1K [82] Structured [8]
@@
Tables 9, 10, and 11 provide per-task results for MAE [21], as summarized in Table 2. Similarly, Tables 12, 13, and 14 present per-task results for MoCo v3 [4], which are also summarized in Table 2.
@@ E.3 Statistical Significance Tests
$p$-value 5.7e-06 1.9e-06 1.9e-06 1.9e-06 0.0001
To determine whether VAPT consistently surpasses competing PEFT methods across the 19 VTAB tasks, we conducted a one-tailed paired Wilcoxon signed-rank test [74] for each VAPT-baseline comparison. The null hypothesis $H_{0}$ for each comparison was that the median difference in performance scores between VAPT and the baseline method is zero. The alternative hypothesis $H_{1}$ was that VAPT performs better than the baseline method. As reported in Table 15, all resulting $p$-values were below the 0.05 significance level, indicating that the observed performance improvements of VAPT are statistically significant.
@@ E.4 Different Backbone Scales
In Figure 4, we compare VAPT and VPT on the CUB-200-2011 dataset [72] using pre-trained backbones of varying scales (ViT-Base, ViT-Large, and ViT-Huge). The results demonstrate that VAPT consistently surpasses VPT as the model size increases. Notably, with the ViT-Huge backbone, VAPT achieves up to a 1.3% improvement over VPT. These findings highlight the scalability and effectiveness of VAPT compared to VPT as model size grows.
@@ E.5 Semantic Segmentation
Table 16: Semantic segmentation results on ADE20K. All methods are evaluated with SETR [84] using ViT-L. The best mIoU scores among all methods but Full are bolded.
@@ E.6 Ablation Study
Components Tuned VTAB-1K [82] Mean Total
@@
Impact of Different Components. We conducted ablation studies on the VTAB-1K benchmark [82] to evaluate the individual contributions of each component in VAPT. The results, presented in Table 17, show that incorporating the channel-wise convolution layer increases the overall average performance by 0.97%. This improvement highlights the benefit of explicitly encoding spatial relationships in the feature map before it is processed by the token-wise projectors. Furthermore, the channel-wise convolution layer reduces the overall parameter count by downsampling the feature map from $H\times W$ to $H^{\prime}\times W^{\prime}$, where $H^{\prime}=H-K+1$ and $W^{\prime}=W-K+1$. Consequently, the dimensionality for each token-wise projector is also reduced, leading to fewer parameters. When the feature projector is removed, performance decreases; for instance, in VTAB-1K Natural, the performance drops from 81.43% to 79.43%. This decrease highlights the importance of the feature projector. Furthermore, our sharing mechanism for the feature projector not only reduces the number of parameters but also facilitates knowledge transfer between layers, leading to improved performance. Overall, combining all components yielded the best performance, with an average accuracy of 72.91% on VTAB-1K.
@@
Method Tuned VTAB-1K [82] Mean Total
@@
Detailed Analysis of the Channel-wise Convolution Layer. We evaluated the effectiveness of our channel-wise convolution layer by comparing its performance to that of a standard convolution layer. We also explored alternative strategies for modeling spatial relationships, including an average pooling layer [41]. Notably, average pooling can be considered a special case of our channel-wise convolution layer, where the kernel weights are fixed rather than learned. The comparative results are presented in Table 18. Our channel-wise convolution layer not only reduces the number of parameters compared to a standard convolution layer but also mitigates overfitting, thereby improving overall performance. Furthermore, it surpasses the average pooling layer; for instance, on VTAB-1K Structured, we achieve a 0.93% performance gain. This improvement is attributed to the greater flexibility of our channel-wise convolution layer compared to that of average pooling.
@@
Method Tuned VTAB-1K [82] Mean Total

### v6

@@ E.1 Per-task Results for VTAB-1K and FGVC
Table 5: VTAB-1K Natural per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in $\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
ViT-B/16 (Dosovitskiy, 2020) VTAB-1K (Zhai et al., 2019) Natural [7]
@@
Full (Iofinova et al., 2022) 68.9 87.7 64.3 97.2 86.9 87.4 38.8 75.88
Linear (Iofinova et al., 2022) 63.4 85.0 63.2 97.0 86.3 36.6 51.0 68.93 [1]
Partial-1 (Yosinski et al., 2014) 66.8 85.9 62.5 97.3 85.5 37.6 50.6 69.44 [2]
MLP-2 (Chen et al., 2020) 63.2 84.8 60.5 97.6 85.9 34.1 47.8 67.70 [2]
MLP-3 (Chen et al., 2020) 63.8 84.7 62.3 97.4 84.7 32.5 49.2 67.80 [2]
MLP-5 (Chen et al., 2020) 59.3 84.4 59.9 96.1 84.4 30.9 46.8 65.98 [1]
MLP-9 (Chen et al., 2020) 53.1 80.5 53.9 95.1 82.6 24.4 43.7 61.90 [1]
Sidetune (Zhang et al., 2020) 60.7 60.8 53.6 95.5 66.7 34.9 35.3 58.21 [0]
Bias (Rebuffi et al., 2017) 72.8 87.0 59.2 97.5 85.3 59.9 51.4 73.30 [3]
Adapter-256 (Cai et al., 2020) 74.1 86.1 63.2 97.7 87.0 34.6 50.8 70.50 [4]
Adapter-64 (Cai et al., 2020) 74.2 85.8 62.7 97.6 87.2 36.3 50.9 70.65 [4]
Adapter-8 (Cai et al., 2020) 74.2 85.7 62.7 97.8 87.2 36.4 50.7 70.67 [4]
VPT-Shallow (Jia et al., 2022) 77.7 86.9 62.6 97.5 87.3 74.5 51.2 76.81 [4]
@@
Table 6: VTAB-1K Specialized per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in $\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
ViT-B/16 (Dosovitskiy, 2020) VTAB-1K (Zhai et al., 2019) Specialized [4]
@@
Full (Iofinova et al., 2022) 79.7 95.7 84.2 73.9 83.36
Linear (Iofinova et al., 2022) 78.5 87.5 68.6 74.0 77.16 [1]
Partial-1 (Yosinski et al., 2014) 78.6 89.8 72.5 73.3 78.53 [0]
MLP-2 (Chen et al., 2020) 74.3 88.8 67.1 73.2 75.86 [0]
MLP-3 (Chen et al., 2020) 77.0 88.0 70.2 56.1 72.83 [0]
MLP-5 (Chen et al., 2020) 73.7 87.2 64.8 71.5 74.31 [0]
MLP-9 (Chen et al., 2020) 78.5 83.0 60.2 72.3 73.49 [0]
Sidetune (Zhang et al., 2020) 58.5 87.7 65.2 61.0 68.12 [0]
Bias (Rebuffi et al., 2017) 78.7 91.6 72.9 69.8 78.25 [0]
Adapter-256 (Cai et al., 2020) 76.3 88.0 73.1 70.5 76.98 [0]
Adapter-64 (Cai et al., 2020) 76.3 87.5 73.7 70.9 77.10 [0]
Adapter-8 (Cai et al., 2020) 76.9 89.2 73.5 71.6 77.80 [0]
VPT-Shallow (Jia et al., 2022) 78.2 92.0 75.6 72.9 79.66 [0]
@@
Table 7: VTAB-1K Structured per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in $\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
ViT-B/16 (Dosovitskiy, 2020) VTAB-1K (Zhai et al., 2019) Structured [8]
@@
Full (Iofinova et al., 2022) 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1 47.64
Linear (Iofinova et al., 2022) 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2 26.84 [0]
Partial-1 (Yosinski et al., 2014) 41.5 34.3 33.9 61.0 31.3 32.8 16.3 22.4 34.17 [0]
MLP-2 (Chen et al., 2020) 45.2 31.6 31.8 55.7 30.9 24.6 16.6 23.3 32.47 [0]
MLP-3 (Chen et al., 2020) 47.8 32.8 32.3 58.1 12.9 21.2 15.2 24.8 30.62 [0]
MLP-5 (Chen et al., 2020) 50.8 32.3 31.5 56.4 7.5 20.8 14.4 20.4 29.23 [0]
MLP-9 (Chen et al., 2020) 47.5 27.9 28.9 54.0 6.2 17.7 10.8 16.2 26.15 [0]
Sidetune (Zhang et al., 2020) 27.6 22.6 31.3 51.7 8.2 14.4 9.8 21.8 23.41 [0]
Bias (Rebuffi et al., 2017) 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1 44.09 [2]
Adapter-256 (Cai et al., 2020) 45.7 37.4 31.2 53.2 30.3 25.4 13.8 22.1 32.39 [0]
Adapter-64 (Cai et al., 2020) 42.9 39.9 30.4 54.5 31.9 25.6 13.5 21.4 32.51 [0]
Adapter-8 (Cai et al., 2020) 45.2 41.8 31.1 56.4 30.4 24.6 13.2 22.0 33.09 [0]
VPT-Shallow (Jia et al., 2022) 50.5 58.6 40.5 67.1 68.7 36.1 20.2 34.1 46.98 [4]
@@
Table 8: FGVC per-task results for ViT-B/16 supervised pre-trained on ImageNet-21K. “Number of Wins” in $\cdot$] denotes the number of wins relative to full fine-tuning. “Tuned/Total” represents the percentage of parameters tuned for each task. The highest accuracy among all approaches except Full is highlighted in bold.
ViT-B/16 (Dosovitskiy, 2020) FGVC [5]
@@
Full (Iofinova et al., 2022) 87.3 82.7 98.8 89.4 84.5 88.54
Linear (Iofinova et al., 2022) 85.3 75.9 97.9 86.2 51.3 79.32 [0]
Partial-1 (Yosinski et al., 2014) 85.6 77.8 98.2 85.5 66.2 82.63 [0]
MLP-2 (Chen et al., 2020) 85.7 77.2 98.2 85.4 54.9 80.28 [0]
MLP-3 (Chen et al., 2020) 85.1 77.3 97.9 84.9 53.8 79.80 [0]
MLP-5 (Chen et al., 2020) 84.2 76.7 97.6 84.8 50.2 78.71 [0]
MLP-9 (Chen et al., 2020) 83.2 76.0 96.2 83.7 47.6 77.31 [0]
Sidetune (Zhang et al., 2020) 84.7 75.8 96.9 85.8 48.6 78.35 [0]
Bias (Rebuffi et al., 2017) 88.4 84.2 98.8 91.2 79.4 88.41 [3]
Adapter-256 (Cai et al., 2020) 87.2 84.3 98.5 89.9 68.6 85.70 [2]
Adapter-64 (Cai et al., 2020) 87.1 84.3 98.5 89.8 68.6 85.67 [2]
Adapter-8 (Cai et al., 2020) 87.3 84.3 98.4 88.8 68.4 85.46 [1]
VPT-Shallow (Jia et al., 2022) 86.7 78.8 98.4 90.7 68.7 84.62 [1]
@@ E.2 Per-task Results on MAE and MoCo v3
Table 9: VTAB-1K Natural per-task results for ViT-B/16 pre-trained on MAE (He et al., 2022). “Number of Wins” in $\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-B/16 (Dosovitskiy, 2020) VTAB-1K (Zhai et al., 2019) Natural [7]
@@
Table 10: VTAB-1K Specialized Per-Task Results for ViT-B/16 Pre-trained on MAE (He et al., 2022). “Number of Wins” in $\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-B/16 (Dosovitskiy, 2020) VTAB-1K (Zhai et al., 2019) Specialized [4]
@@
Table 11: VTAB-1K Structured Per-Task Results for ViT-B/16 Pre-trained on MAE (He et al., 2022). “Number of Wins” in $\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-Base/16 (Dosovitskiy, 2020) VTAB-1K (Zhai et al., 2019) Structured [8]
@@
Table 12: VTAB-1K Natural Per-Task Results for ViT-B/16 Pre-trained on MoCo v3 (Chen et al., 2021). “Number of Wins” in $\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-B/16 (Dosovitskiy, 2020) VTAB-1K (Zhai et al., 2019) Natural [7]
@@
Table 13: VTAB-1K Specialized Per-Task Results for ViT-B/16 Pre-trained on MoCo v3 (Chen et al., 2021). “Number of Wins” in $\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-B/16 (Dosovitskiy, 2020) VTAB-1K (Zhai et al., 2019) Specialized [4]
@@
Table 14: VTAB-1K Structured Per-Task Results for ViT-B/16 Pre-trained on MoCo v3 (Chen et al., 2021). “Number of Wins” in $\cdot$] denotes comparisons relative to full fine-tuning. “Tuned/Total” is the percentage of parameters tuned for each task. The highest accuracy among all approaches is highlighted in bold.
ViT-Base/16 (Dosovitskiy, 2020) VTAB-1K (Zhai et al., 2019) Structured [8]
@@
Tables 9, 10, and 11 provide per-task results for MAE (He et al., 2022), as summarized in Table 2. Similarly, Tables 12, 13, and 14 present per-task results for MoCo v3 (Chen et al., 2021), which are also summarized in Table 2.
@@ E.3 Statistical Significance Tests
pp-value 5.7e-06 1.9e-06 1.9e-06 1.9e-06 0.0001
To determine whether VAPT consistently surpasses competing PEFT methods across the 19 VTAB tasks, we conducted a one-tailed paired Wilcoxon signed-rank test (Wilcoxon, 1992) for each VAPT-baseline comparison. The null hypothesis $H_{0}$ for each comparison was that the median difference in performance scores between VAPT and the baseline method is zero. The alternative hypothesis $H_{1}$ was that VAPT performs better than the baseline method. As reported in Table 15, all resulting $p$-values were below the 0.05 significance level, indicating that the observed performance improvements of VAPT are statistically significant.
@@ E.4 Different Backbone Scales
In Figure 4, we compare VAPT and VPT on the CUB-200-2011 dataset (Wah et al., 2011) using pre-trained backbones of varying scales (ViT-Base, ViT-Large, and ViT-Huge). The results demonstrate that VAPT consistently surpasses VPT as the model size increases. Notably, with the ViT-Huge backbone, VAPT achieves up to a 1.3% improvement over VPT. These findings highlight the scalability and effectiveness of VAPT compared to VPT as model size grows.
@@ E.5 Semantic Segmentation
Table 16: Semantic segmentation results on ADE20K. All methods are evaluated with SETR (Zheng et al., 2021) using ViT‑L. The best mIoU scores among all methods but Full are bolded.
@@ E.6 Ablation Study
Components Tuned VTAB-1K (Zhai et al., 2019) Mean Total
@@
Impact of Different Components. We conducted ablation studies on the VTAB-1K benchmark (Zhai et al., 2019) to evaluate the individual contributions of each component in VAPT. The results, presented in Table 17, show that incorporating the channel-wise convolution layer increases the overall average performance by 0.97%. This improvement highlights the benefit of explicitly encoding spatial relationships in the feature map before it is processed by the token-wise projectors. Furthermore, the channel-wise convolution layer reduces the overall parameter count by downsampling the feature map from $H\times W$ to $H^{\prime}\times W^{\prime}$, where $H^{\prime}=H-K+1$ and $W^{\prime}=W-K+1$. Consequently, the dimensionality for each token-wise projector is also reduced, leading to fewer parameters. When the feature projector is removed, performance decreases; for instance, in VTAB-1K Natural, the performance drops from 81.43% to 79.43%. This decrease highlights the importance of the feature projector. Furthermore, our sharing mechanism for the feature projector not only reduces the number of parameters but also facilitates knowledge transfer between layers, leading to improved performance. Overall, combining all components yielded the best performance, with an average accuracy of 72.91% on VTAB-1K.
@@
Method Tuned VTAB-1K (Zhai et al., 2019) Mean Total
@@
Detailed Analysis of the Channel-wise Convolution Layer. We evaluated the effectiveness of our channel-wise convolution layer by comparing its performance to that of a standard convolution layer. We also explored alternative strategies for modeling spatial relationships, including an average pooling layer (LeCun et al., 1998). Notably, average pooling can be considered a special case of our channel-wise convolution layer, where the kernel weights are fixed rather than learned. The comparative results are presented in Table 18. Our channel-wise convolution layer not only reduces the number of parameters compared to a standard convolution layer but also mitigates overfitting, thereby improving overall performance. Furthermore, it surpasses the average pooling layer; for instance, on VTAB-1K Structured, we achieve a 0.93% performance gain. This improvement is attributed to the greater flexibility of our channel-wise convolution layer compared to that of average pooling.
@@
Method Tuned VTAB-1K (Zhai et al., 2019) Mean Total